An object of mass 10kg on a flat horizontal surface is pulled with a force of 70N at an angle of 30 degrees above the horizontal. The object slides along the surface with an acceleration coefficient of k.e is 0.3. What is the normal force.
1
Expert's answer
2020-05-07T09:49:16-0400

The normal force


"N=mg-F\sin30°=10\cdot 9.81-70\cdot \sin30°=63.1 N"